The finance manager of ABC ltd has computed the after tax marginal costs of the company;s three sources of capital ordinary shares-600000 at a cost of 14%600000-1500000 at a cost of 17%over 1500000 at cost of 18%preference shares0-300000 at a cost of 12%over 300000 at a cost of 13%debts0-1000000 at a cost of 5%over 1000000 at a cost of 8%the manager wants to maintain the current present capital structure of 50% debt , 10% preference shares and 40% ordinary shares.Requiredi) the break point in the capital structure (ii) the weighted average cost of capital between breakpoints iii) draw the weighted marginal cost of capital schedule
Answer #1
ordinary shares Cost of equity prefference shares cost of Pref shares Debt after tax cost of debt
0-600,000 14% 0-300,000 12% 0-1,000,000 5%
600,000-1,500,000 17% 300000 above 13% abv 1,000,000 8%
above 1,500,000 18%
capital structure 40% 10% 50%
1 the break point is:
break point is the point whre the capital amount of capital changes as a structre
here it can be seen that it happens so in first case
the break point is: (600000+300000+100000)= $                 1,900,000
2 WACC= {kd (1-t)*debt/ debt+ equity}+ {ke*equity/debt+ equity}
WACC 1=(0.14*0.40)+(0.12*0.10)+(0.05*0.50) 9.30%
WACC 2= (0.17*0.40)+(0.13*0.10)+(0.08*0.50) 12.10%
it can be seen that the company can invest upto $1900000 at 9.30%
amd if it increases its capital, the WACC will increase to 12.10%

3.